Describe (list the elements, give the identity and inverses) the cyclic


group generated (under multiplication) by [ 1 1 ]

[0 1] matrix



1
Expert's answer
2021-02-24T07:48:46-0500

Solution:

Let given matrix be "A=\begin{bmatrix}1&1\\ 0&1\end{bmatrix}"

Now, "A^2=\begin{bmatrix}1&1\\ 0&1\end{bmatrix}^2=\begin{bmatrix}1&2\\ 0&1\end{bmatrix}"

Next, "A^3=A^2.A=\begin{bmatrix}1&2\\ 0&1\end{bmatrix}.\begin{bmatrix}1&1\\ 0&1\end{bmatrix}=\begin{bmatrix}1&3\\ 0&1\end{bmatrix}"

And so on.

Then, we observe that "A^n=A^{n-1}.A=\begin{bmatrix}1&n-1\\ 0&1\end{bmatrix}.\begin{bmatrix}1&1\\ 0&1\end{bmatrix}=\begin{bmatrix}1&n\\ 0&1\end{bmatrix}\ne\begin{bmatrix}1&0\\ 0&1\end{bmatrix}"

We see that for no value of "n" , we have "A^n=I" .

Thus, given matrix A is not cyclic.